OpenStudy (anonymous):

What is the integral of (x+(1/x))^2 on [1,2]?

OpenStudy (anonymous):

i would square it first, then compute the integral

OpenStudy (anonymous):

in other words, compute \[\int_0^1x^2+2+\frac{1}{x^2}dx\] break in to three parts, the last on is an improper integral

OpenStudy (anonymous):

oh no it isn't sorry, the limits are from 1 to 2 compute \[\int_1^2\left(x^2+2+\frac{1}{x^2}\right)dx\]

OpenStudy (anonymous):

So would I subtract the antiderivative of 1 from the antiderivative of 2.

OpenStudy (anonymous):

NEVERMIND! Thanks!